OpenStudy (anonymous):

Find in terms of the positive integer k, the sum of the integers from 2k to 4k inclusively

OpenStudy (zarkon):

just use the formula (twice) \[\sum_{x=1}^{n}x=\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

how did you know that is the formula to use?

OpenStudy (anonymous):

think of it like this

OpenStudy (anonymous):

how do you know the conjecture is 'i'?

OpenStudy (anonymous):

Whenever there is a question about adding up consecutive integers (like 1+2+3+4+...+15, or 22+23+24+...+50) that formula should be the first thing that pops in your mind.

OpenStudy (anonymous):

the 'i' is just a dummy variable, a place holder. If you look at zarkon's formula, he used x instead. It means the same thing. we could pick any letter we want for that purpose :)

OpenStudy (anonymous):

but the 'i ' beside the summation symbol what does it mean?

OpenStudy (zarkon):

\[\sum_{\Omega=1}^{\Gamma}\Omega=\frac{\Gamma(\Gamma+1)}{2}\] ;)

OpenStudy (anonymous):

@ zarkon lol that makes it look awesome The 'i' next to the summation is the formula (or function) you are plugging each value of the index into. For example: \[\sum_{i=1}^{5}i = 1+2+3+4+5\] or \[\sum_{i=5}^{10}i^2 = 5^2+6^2+7^2+8^2+9^2+10^2\] In general: \[\sum_{i=a}^{b}f(i) = f(a)+f(a+1)+f(a+2)+\cdots +f(b-1)+f(b)\]

OpenStudy (anonymous):

i hope thats not too confusing >.< im sry if it is.

OpenStudy (anonymous):

how do you know that is the formula to use? it only tells me from 2k to 4k which really doesn't tell me much..

OpenStudy (anonymous):

Well, you are adding up all the integers in between 2k and 4k, so in my mind im thinking: \[2k+(2k+1)+(2k+2)+(2k+3)+\cdots+(4k-2)+(4k-1)+4k\]

OpenStudy (anonymous):

if k was equal to say...3, then we would be adding the numbers from 6 to 12: \[6+7+8+9+10+11+12\]

OpenStudy (anonymous):

so i can say it's like an adding up series and that is why we use that formula (1/2)(n)(n+1) which is like an AP summation?

OpenStudy (anonymous):

right right :)

OpenStudy (anonymous):

hmmm i think i get it.. i shall try it out =)

OpenStudy (anonymous):

joemath! now the question asks for the sum of all odd integers lying between 2k and 4k what do i do???

OpenStudy (anonymous):

hmm...we need a formula for the sum of the odd integers...unfortunately i dont know one off the top of my head, maybe Zarkon does? i'll have to work out one on paper. one sec.

OpenStudy (zarkon):

\[3k^2\]

OpenStudy (anonymous):

Alright i came up with a formula for the sum of all the odd numbers up to a given odd number (say, n). its: \[1+3+5+7+\cdots+(n-2)+n = \frac{(n+1)^2}{4}\] Note that n must be odd. For example: \[1+3+5+7 = \frac{(7+1)^2}{4} = \frac{8^2}{4} = \frac{64}{4} = 16\]

OpenStudy (zarkon):

\[\sum_{i=k}^{2k-1}{2i+1}\] first term is 2k+1 and last term is 2(2k-1)+1=4k-2+1=4k-1. and it is counting the odd numbers

OpenStudy (anonymous):

thing is we must have a fromula for summing up odd numbers?

OpenStudy (zarkon):

\[\sum_{i=k}^{2k-1}{2i+1}=3k^2\]

OpenStudy (anonymous):

ah, that what we need, nice Zarkon.

OpenStudy (zarkon):

sorry it took so long...I got distracted by the dead bird my wife found in the back yard :(

OpenStudy (anonymous):

first odd term is 2k+1?

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

yes, 2k is even (because its divisible by 2). So 2k+1 has to be the first odd term.

OpenStudy (anonymous):

ok... so we have to sum up from 2k to 4k odd terms... first term is 2k+1... then how do i proceed??

OpenStudy (anonymous):

That summation zarkon posted works out perfectly: \[\sum_{i=k}^{2k-1}2i+1 = (2k+1)+(2(k+1)+1)+(2(k+2)+1)+\cdots + (2(2k-1)+1)\] =\[(2k+1)+(2k+3)+(2k+5)+\cdots +(4k-1)\] It hits all the odd numbers. So we just have to figure out the sum, which i'll do on paper and post.

OpenStudy (anonymous):

sigh.. i don't understand..

OpenStudy (zarkon):

can you be more specific about what you don't understand?